A disk starts spinning from rest at t=0, with a constant angular
acceleration of 1.2 rad/s2. The radius of the disk is 0.4 m.
Consider a bug clinging to the outer edge of the disk. What is the
magnitude of the acceleration felt by this bug at t=1 s? Choose the
closest answer.
Answer #1

w = w0 + alpha*t

w0 = 0

w = 1.2*1

=1.2

centreletel acceleration = w2r

= 1.2*1.2*0.4

= 0.576

tangential acc = r*alpha = 1.2*0.4 = 0.48

total acc = sqrt(alpha2+centreptel2)

=sqrt(0.48*0.48+0.576*0.576)

= 0.749